OpenStudy (anonymous):

F(x) = 8x^3 - 3x^2 + 5x +15

OpenStudy (unklerhaukus):

\[F(x) = 8x^3 - 3x^2 + 5x +15\\[2ex] f(x)=\frac{\mathrm d}{\mathrm dx}F(x)\\ \qquad=\frac{\mathrm d}{\mathrm dx}\Big(8x^3 - 3x^2 + 5x +15\Big)\\ \qquad=\frac{\mathrm d}{\mathrm dx}\Big(8x^3\Big) - \frac{\mathrm d}{\mathrm dx}\Big(3x^2 \Big)+ \frac{\mathrm d}{\mathrm dx}\Big(5x\Big) +\frac{\mathrm d}{\mathrm dx}\Big(15\Big)\\[3ex] \qquad\hspace{6em}\textit{< applying }\quad \frac{\mathrm d}{\mathrm dx}ax^n=nax^{n-1} \textit{ >}\\[3ex] \qquad=3\cdot8x^{3-1} - 2\cdot3x^{2-1} + 1\cdot5x^{1-1} +0\cdot15x^{0-1}\\ \qquad= \]

OpenStudy (anonymous):

What's the answer

OpenStudy (unklerhaukus):

What's the question?